OpenStudy (anikate):

divide the rtional expressions. Simplify the result. (48x2/y)÷(36xy2/5)

OpenStudy (diyadiya):

When you divide 2 fractions, you multiply the first fraction with the reciprocal of the second \[\frac{48x^2}{y} \times \frac{5}{36xy^2} \]

OpenStudy (anonymous):

\[\huge \color{green} {\frac {\frac{a}{b}}{\frac{c}{d}}} = \color{blue}{\frac{a \times d}{b \times c}}\]

OpenStudy (jiteshmeghwal9):

sorry ! there is y^3

OpenStudy (anonymous):

@jiteshmeghwal9 do you still think your answer is right??

OpenStudy (anikate):

yup

OpenStudy (anikate):

one more

OpenStudy (anonymous):

I guess no..

OpenStudy (jiteshmeghwal9):

there is x^3 also

OpenStudy (jiteshmeghwal9):

at numerator

OpenStudy (anonymous):

Again @jiteshmeghwal9 do you still think you answer is right???

OpenStudy (anonymous):

Are you sure about \(\large x^3\) in the numerator?? @jiteshmeghwal9

OpenStudy (diyadiya):

@Anikate Can you show your work? What did you get?

OpenStudy (anikate):

\[(x ^{2}/x ^{2}-1)/(3x/x+1)\]

OpenStudy (anonymous):

Is this the another question??

OpenStudy (jiteshmeghwal9):